OpenStudy (anonymous):

Calculas help? ∫??(5−ex)(e2x)dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?}(5-e ^{x})(e ^{2x})\]

OpenStudy (bahrom7893):

Multiply it out, it's gonna make your life much easier

OpenStudy (bahrom7893):

Int (5e^(2x)-e^(3x),x)

OpenStudy (anonymous):

opps it is suppose to be divided by the second e

OpenStudy (bahrom7893):

ohh that makes it a little different

OpenStudy (anonymous):

yea i would think so lol

OpenStudy (bahrom7893):

oh still rewrite it as: Int (5/e^(2x) - e^x/e^(2x))) Int (5*e^(-2x) - e^(-x))

OpenStudy (bahrom7893):

Now split it into two: 5 * Int(e^(-2x)) - Int(e^(-x))

OpenStudy (bahrom7893):

Now just use the rule: Int (e^(kx)) = e^(kx) / k

OpenStudy (bahrom7893):

In the first half k = -2, in the second half, k = -1 So finally: 5 * (e^(-2x)/(-2)) + e^(-x) + C

OpenStudy (anonymous):

ok that makes sense thanks

OpenStudy (bahrom7893):

btw to prove that Int (e^(kx)) = e^(kx)/k just use u sub. Let u = kx, du = k dx, etc..